### v1

# AI-Enhanced Self-Triggering for Extensive Air Showers: Performance and FPGA Feasibility

Abstract
Cosmic-ray detection with radio antennas has traditionally depended on external triggers from particle detectors, constraining sensitivity and increasing complexity. Previous attempts at fully standalone, radio-only triggers have often failed under intense radio frequency interference, making genuine air-shower signals difficult to isolate. We present a proof-of-principle artificial intelligence-based self-triggering system that overcomes these limitations. By training a deep learning model on both real noise data and injected cosmic-ray-like pulses, we achieve an exceptionally low false-positive rate alongside high detection efficiency. Configurable operating points can suppress false positives below 0.01% while retaining more than 88% of genuine signals, and can even eliminate false positives entirely at a modest reduction in signal efficiency. This flexibility makes single-station cosmic-ray detection feasible without requiring external trigger inputs. Applying our approach to real-world noise conditions reduces the initial false-positive event rate by several orders of magnitude, supporting large-scale deployments. Extrapolation to dedicated hardware implementations, such as FPGAs, indicates that sub- inference times are achievable, enabling real-time autonomous triggering. These results highlight the transformative potential of artificial intelligence for enhancing radio detection sensitivity and inaugurate a new generation of fully self-triggered cosmic-ray observatories.
1 Introduction
Ultra-high-energy cosmic rays (UHECRs) are primarily detected using large-scale ground-based observatories that employ a combination of particle detectors, fluorescence telescopes, and radio antennas [1]. Each technique provides complementary insights into extensive air showers (EAS), enabling detailed studies of cosmic-ray composition, energy spectrum, and arrival direction. Among these methods, radio detection has gained increasing attention due to its cost-effectiveness, full-sky coverage, and near-continuous operation [2]. By measuring the radio emission generated primarily by geomagnetic deflection [3] and the charge-excess effect [4] in the EAS, radio arrays can reconstruct key shower parameters, such as the atmospheric depth of the shower maximum () and primary energy, with high precision [5, 6].
A distinct advantage of radio detection is its sensitivity to inclined air showers, where traditional particle detector arrays struggle due to the attenuation of the electromagnetic shower component in the atmosphere [6]. However, despite these benefits, current radio detection techniques rely on external triggers from other detector systems, such as surface or fluorescence detectors, limiting their autonomy and increasing operational complexity. The development of an effective self-triggering system is therefore crucial to fully exploit the potential of radio detection for cosmic-ray observations [7].
Previous attempts to implement self-triggering in radio arrays, such as those by CODALEMA and AERA, relied on conventional techniques like narrowband radio frequency interference (RFI) suppression and threshold-based triggering [8, 9]. However, these methods proved ineffective in distinguishing cosmic-ray-induced radio pulses from transient noise sources, leading to a high rate of false positives. As a result, these efforts faced significant challenges and did not produce established scientific results, underscoring the need for more advanced signal-processing strategies.
Recent developments in modelling EAS radio emissions, combined with experimental and simulation studies, have provided a deeper understanding of characteristic pulse shapes and ground radiation patterns [2]. This progress has paved the way for more sophisticated signal-processing techniques, including artificial intelligence (AI)-driven approaches for enhancing radio-based self-triggering. Recent studies show that deep learning models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), can effectively distinguish cosmic-ray signals from background noise, significantly reducing detection thresholds. These techniques have been successfully applied to identify air-shower radio pulses in offline analyses [10], demonstrating their potential to enhance signal selection. Recent efforts, such as those in the GRAND experiment, continue to investigate self-triggering using AI-based methods, including CNNs. While preliminary results indicate that these methods can reject at least 40% of background noise at a 90% signal selection efficiency [11], achieving real-time deployment remains challenging. Self-triggering requires ultra-low latency processing to ensure the identification of cosmic-ray signals within microseconds of their arrival, while also maintaining a false-positive trigger rate at the sub- level, a constraint that traditional software-based neural networks struggle to meet.
To address the challenge of self-triggering in cosmic-ray radio signal detection, AI-accelerated signal processing implemented on Field-Programmable Gate Arrays (FPGAs) offers a viable solution. FPGAs provide exceptional parallel processing capabilities and low-latency inference, making them ideal for real-time feature extraction and classification of radio pulses. Similar approaches have been explored in high-energy physics experiments, such as the ATLAS Tile Calorimeter, where deep learning algorithms implemented on FPGAs have demonstrated significant improvements in real-time signal reconstruction under high pileup conditions [12, 13]. These advancements highlight the potential of FPGA-based AI techniques in handling complex, high-rate signal environments, which are also crucial for enhancing the efficiency and accuracy of cosmic-ray self-triggering.
In this work, we investigate the feasibility of FPGA-based AI self-triggering for radio cosmic-ray detection. We begin by training and evaluating our deep neural network on Metal Performance Shaders (MPS) with data that includes experimentally measured background noise, then estimate their potential performance on FPGAs. While high-level synthesis (HLS) frameworks such as HLS4ML [14] and Xilinx Vitis AI [15] offer streamlined approaches to deploying AI models on FPGAs, we consider their integration a logical extension of this proof-of-concept study rather than part of our immediate scope.
Our method comprises a fully convolutional autoencoder that denoises radio traces by suppressing background noise while retaining cosmic-ray pulses, followed by a fully convolutional classifier (FCN) that differentiates cosmic-ray signals from noise. By leveraging real data with measured background noise, we enhance the robustness of our approach in practical deployments. We also address critical issues associated with FPGA-based AI, including quantization strategies, resource utilization, and latency reduction, and propose a framework for real-time classification of cosmic-ray radio pulses. This proof-of-principle demonstration paves the way for future work incorporating more advanced HLS techniques to further optimize deployment on FPGA hardware.
The implementation of AI-accelerated FPGA-based self-triggering represents a significant advancement in radio detection technology. By reducing reliance on external detector triggers and increasing detection efficiency, this approach enables greater autonomy for radio observatories, facilitating the detection of rare and high-energy events, including ultra-high-energy neutrinos. Our results demonstrate the feasibility of deploying AI-driven self-triggering at the detector level, marking an important step toward fully autonomous, real-time radio detection systems for astroparticle physics.
2 Experimental Setup
The dataset used in this study consists of experimentally recorded noise traces collected in the physics campus of the University of Siegen. The measurements were performed using a butterfly antenna connected to a low-noise amplifier with 30 dB gain, which was further linked to a digital oscilloscope. These recordings capture real-world transient noise sources, including atmospheric and anthropogenic contributions.
To focus on the frequency range relevant for air shower radio detection, the recorded signals were digitally bandpass-filtered between 30 and 80 MHz. Furthermore, data was sampled at to ensure comparability with real-world cosmic-ray radio detectors, such as AERA and Auger RD, which operate at and , respectively. This adjustment reflects the practical constraints of large-scale detectors, where lower sampling rates are used due to budget and hardware limitations, ensuring the applicability of the method to future radio arrays.
The dataset used in this study was collected in an urban environment, where background noise levels are higher and more variable compared to remote cosmic-ray observatories. To capture a diverse range of real-world noise conditions, data was collected over approximately two hours, yielding 200 million samples. Data was continuously streamed, with traces read out at a rate of one per second. Given the high variability of the urban noise environment, the dataset presents a challenging test case for AI-driven signal detection methods. This setting introduces additional challenges, including increased transient interference and strong radio frequency interference (RFI). Two prominent narrowband RFI sources were identified at approximately and , along with other narrowband and transient RFI. The power spectrum of the recorded traces, shown in Figure 1, was obtained after applying a digital bandpass filter in the range of 30 to . For the initial analysis, these narrowband RFIs were not suppressed to assess the model’s robustness in the presence of strong interference. The results presented in the following section reflect model performance when both training and testing were conducted on data sets containing these RFIs. Additionally, we examine how performance changes when these narrowband RFIs are mitigated prior to training and validation, providing insights into the impact of RFI suppression on the learning process, which we will report in the discussion section 7.
To evaluate the AI model’s performance, signal-like pulses were injected into measured noise traces at random positions, covering a broad range of amplitudes. The pulse power was scaled relative to the noise power, which was calculated using a pre-filtered version of each trace. This ensured that the injected pulses maintained a well-controlled signal-to-noise ratio (SNR) within a user-defined range while accurately reflecting real-world noise conditions. However, to prevent double filtering of the noise, which could distort its characteristics, the pulses were injected into the original (unfiltered) traces. This step also eliminated systematic biases by ensuring that the statistical properties of the background noise remained unaffected by the injection process. To maintain consistency in data processing, the digital bandpass filter was applied only after pulse injection, ensuring that both noise-only and signal-injected datasets underwent identical filtering and conditioning. This approach was chosen over COREAS simulations [16] to establish a proof of principle based purely on signal strength, avoiding dependencies on the absolute energy scale, antenna gain pattern, and shower inclination, which introduce additional uncertainties. By injecting short pulses into measured noise traces from a real-world, high-interference environment, we enable a focused evaluation of the AI model’s ability to detect transient signal shapes under conditions where traditional threshold-based triggering systems struggle or fail entirely due to overwhelming radio frequency interference (RFI). This approach enables a systematic assessment of the model sensitivity in varying SNRs while ensuring robustness under challenging conditions. By directly testing and refining AI-based techniques on measured noise traces—without relying on an external event-reconstruction framework or requiring a particle detector—it provides a controlled and scalable test case. Although it does not fully replicate all details of cosmic-ray pulse shapes, much of which may already be washed out by bandpass filtering, it offers a robust estimate of the false positive rate of AI-based triggering methods in a realistic noise environment.
2.1 Training and Validation Datasets
To train and evaluate the model, multiple datasets were used, each designed to fulfil a specific role in ensuring robust performance. The datasets were categorized into training, cross-validation, and independent validation sets to assess the model’s generalization and sensitivity. All noise traces are unique and are derived from experimental measurements, randomly selected from the full duration of the noise measurement campaign, whereas the signal-containing traces consist of injected pulses, as previously described. For the initial analysis, only a bandpass filter was applied, with no additional filtering. However, we later incorporate further filtering to suppress prominent RFI sources in the 30– range for comparison, the results of which are discussed in section 7.
Training and Cross-Validation Datasets: The dataset used for training and model tuning was randomly split into two equal parts:
Training Dataset (50%): Comprised 15,000 time series traces, each containing 128 samples recorded at a sampling rate of , consistent with standard practices in radio-based air shower detection. It included an equal distribution of 5,000 pure noise traces (background), 5,000 traces of isolated EAS-like pulses (pure signal), and 5,000 traces with EAS-like pulses injected into noise (signal). This setup ensured that the model effectively learned to distinguish genuine signals from background noise.
Cross-Validation Dataset (50%): Consisted of another 15,000 traces, also evenly divided into 5,000 background traces, 5,000 pure signal traces, and 5,000 signal traces. This dataset was used to fine-tune the model and monitor its generalization performance during training.
These datasets were randomly derived from a larger pool of available traces, ensuring an unbiased and representative sample distribution.
Independent Validation Datasets: Three independent validation datasets were used to assess the robustness of the model under different conditions.
Validation Dataset 1: Contained 20,000 traces, equally split into 10,000 background and 10,000 signal traces. This dataset was applied to evaluate the performance of the model on unseen data with an SNR distribution similar to the training set.
Validation Dataset 2: Consisted of 20,000 traces, equally divided into 10,000 background traces and 10,000 signal traces with extremely low-amplitude pulses. This dataset was specifically designed to test the sensitivity of the model in low-SNR scenarios.
Validation Dataset 3: Comprised 1 million pure noise traces to assist in evaluating the false positive triggering rate. This dataset provides a large statistical sample to quantify how frequently the model incorrectly classifies noise as a signal event.
These independent validation data sets were crucial to assessing the model’s ability to generalize to previously unseen data and to reliably detect signals under varying noise conditions, particularly at low SNR.
Figure 2 illustrates the SNR distributions for background and signal traces across Validation Dataset 1, which closely follows the distribution of the training dataset, and Validation Dataset 2, which consists of low-amplitude signals. This comparison provides insights into the differences in signal detectability across varying SNR conditions.
3 Methodology
We implemented a deep learning-based pipeline for denoising and classifying radio time series in the context of self-triggering for EAS events. The pipeline consists of two primary components:
1.
A CNN-based denoiser to suppress noise in the input signals.
2.
A fully convolutional classifier (FCN) to distinguish between signal and background in a binary classification task.
These models operate sequentially, where the denoiser first refines raw data before passing it to the classifier for event discrimination.
3.1 Model Architecture
3.1.1 CNN-Based Denoiser
The denoiser is a fully convolutional autoencoder designed to suppress noise and enhance cosmic ray pulses in noisy radio time series.. It follows an encoder-bottleneck-decoder structure with a skip connection to retain fine-grained features.
Encoder: Two convolutional layers extract multi-scale features, followed by a single max pooling layer that reduces resolution by a factor of two.
Bottleneck: A deeper convolutional block refines feature representations at the reduced resolution.
Decoder: An upsampling layer restores the original resolution, followed by convolutional layers that refine the denoised signal. A skip connection concatenates high-resolution features from the encoder to enhance detail recovery.
The denoiser learns from a dataset where the inputs are noisy traces containing EAS-like pulses (signal), and the targets are the corresponding pure EAS-like pulses (pure signal).
The training process minimises the Mean Squared Error (MSE) loss between the clean target and the denoised prediction :
This architecture ensures that the model preserves cosmic ray features while effectively filtering out background noise.
3.1.2 Fully Convolutional Classifier (FCN)
The FCN classifier classifies time series into background or signal, where signal refers to a noisy trace containing an EAS pulse. It includes:
Three convolutional blocks with batch normalization, dropout, and max pooling.
Global average pooling (GAP) to reduce spatial dimensions while retaining features.
A fully connected layer with a sigmoid activation function for binary classification.
The classifier is trained using the Binary Cross-Entropy (BCE) loss:
where is the true label and is the predicted probability.
3.2 Training and Optimization
Both models were trained using the Adam optimizer, with task-specific loss functions, following a two-step training approach:
1.
Denoiser Training:
Trained on traces with pulses injected into noise (signal) and pure signal.
Optimised to minimize Mean Squared Error (MSE) loss, enhancing signal quality while suppressing noise.
2.
Classifier Training:
Trained on denoised signal traces (output of the denoiser) alongside denoised background traces.
Optimized to minimize Binary Cross-Entropy (BCE) loss, ensuring accurate discrimination between signal-containing and noise-only traces.
The model was trained for 15 epochs with a batch size of 64 and a learning rate of 0.001, using the Training Dataset. Its generalization performance was monitored with the Cross-Validation Dataset described in Section 2.1. The model was implemented using PyTorch with support for MPS acceleration. Training was conducted using the Adam optimizer, and model performance was evaluated on the cross-validation dataset using classification accuracy, precision, recall, and F1-score.
4 Model Performance Analysis
The first two plots in Figure 3 (top row) illustrate the probability distributions of sigmoid prediction for validation datasets 1 and 2, which differ in their SNR distributions. In Validation Dataset 1, where the SNR distribution closely resembles that of the training set (see Section 2.1), the model exhibits a strong separation between background and signal traces. Most noise traces receive lower probabilities, while signal-containing traces are predominantly assigned higher values. In contrast, Validation Dataset 2 contains traces with significantly lower SNRs, commonly below the noise floor. While the model still distinguishes between the two classes, the classification performance degrades: the distribution of signal-containing traces is more spread out, with a noticeable tail extending toward lower probabilities. This shift indicates a reduction in signal efficiency, as weaker signals are increasingly misclassified due to their low SNR.
Beyond assessing performance across different SNRs, it is also essential to evaluate the model’s behaviour in environments dominated by overwhelming background noise, where false positives become a critical concern. To investigate this, we introduce Validation Dataset 3, which consists of one million background traces. The bottom row of Figure 3 explores the model’s response to this dataset by comparing background traces from Validation Dataset 3 with signal traces from Validation Dataset 1. The sigmoid probability distribution (bottom left) shows that while most noise traces receive low probabilities, the large dataset size results in a fraction extending into the signal region. This highlights the need to quantify the false trigger rate under realistic high-background conditions. The final plot (bottom right) presents the distributions of the model’s logits, i.e. the raw scores before applying a probability function. These distributions reveal a clear separation between background traces from Validation Dataset 3 and signal traces from Validation Dataset 1. Using logits instead of sigmoid probabilities significantly improves false positive rejection, albeit at the cost of true positive efficiency. This trade-off is particularly relevant when designing self-triggering strategies in noisy environments.
4.1 Performance Metrics and Trigger Rates
Following the qualitative assessment in Figure 3, we systematically evaluate the model’s performance across different validation datasets. This analysis focuses on three key aspects: trigger efficiency, robustness to low-SNR signals, and false trigger rate in high-background conditions. The results are presented as follows.
4.1.1 Trigger Efficiency in Controlled Conditions (Validation Dataset 1)
True Positives: 99.93%
False Positives: 0.07%
A simple probability threshold is sufficient for efficient signal selection in this case.
4.1.2 Robustness to Weaker Pulses (Validation Dataset 2)
True Positives: 87.96%
False Positives: 0.08%
Despite the increased challenge, the model effectively distinguishes signal-containing traces, demonstrating its robustness to weaker pulses. The reduction in efficiency is expected as most of these signals fall outside the training range. Although further tuning could enhance performance, the primary objective here is to assess the model’s ability to generalise rather than optimise for specific low-SNR cases.
4.1.3 False Trigger Rate Estimation in a High-Noise Environment (Validation Dataset 3)
To assess the model’s performance in the presence of overwhelming background noise, we evaluated it on Validation Dataset 3. This dataset simulates realistic high-RFI conditions, such as those found in urban environments (e.g., LOFAR, LOPES) or certain AERA stations [7].
The primary objective of this evaluation is to quantify false trigger rates in such environments. The results are as follows:
True Positives (Initial Threshold): 99.93%
False Positives: 768 per 1M events (0.08%), which translates to approximately 800 Hz first-level trigger (FLT) rate in a 1 MHz trigger scenario.
Applying confidence-based selection criteria (referred here as entropy cut) significantly reduces the false positive rate while maintaining high true positive efficiency.
Entropy Cut: 99.35% True Positives, 79 False Positives per 1M events (0.008%), corresponding to approximately 80 Hz false triggers in a 1 MHz trigger scenario.
The FLT rate is already suitable for a station-level trigger and provides a strong foundation for further refinement. A subsequent second-level trigger (SLT) stage, utilizing array-level coincidence searches, could further suppress false positives while preserving efficiency.
To explore the feasibility of detecting cosmic-ray events at a single station, we further applied logits-based cuts, which resulted in:
Logit Cut: 72.4% True Positives, 0 False Positives, completely eliminating false triggers in this dataset.
This analysis shows that entropy-based selection effectively reduces the trigger rate to an acceptable level for an FLT, while still maintaining high detection efficiency. Further refinement at the SLT stage, through array-level coincidence searches, could optimise false positive rejection even further. On the other hand, applying logits-based cuts enables the possibility of detecting cosmic-ray events at a single station, drastically suppressing false positives, with none observed in our dataset. While this comes with some reduction in signal efficiency, it remains within a reasonable range, making single-station detection feasible in certain scenarios.
4.1.4 Example of a Low-Amplitude Signal Classification
Figure 4 illustrates a representative trace containing a faint signal pulse correctly classified by the model. The pulse is indistinguishable by eye due to transient noise, yet the classifier confidently assigns it a high probability. This result highlights the model’s ability to recover weak air shower signals that would otherwise remain undetectable using traditional threshold-based methods.
5 Estimated Latency on FPGA Platforms
To evaluate the feasibility of real-time deployment, the model’s inference latency was first measured on an Apple Silicon GPU using Metal Performance Shaders (MPS). Each inference took approximately per trace, which implies a required buffer of around 8,250 samples when sampling at . Since real-time streaming at this rate is continuous, reducing inference latency is essential to minimize buffering and avoid potential pipeline stalls.
Moving from a GPU to an FPGA requires a different set of assumptions compared to traditional CPU-to-FPGA speed-up estimates. In particular, FPGAs allow for specialized parallelism and hardware-friendly dataflow optimizations that can significantly reduce latency if designed properly. Table 1 summarizes conservative latency estimates for different FPGA classes, along with the corresponding buffer sizes at . These values were derived by adjusting for GPU-to-FPGA acceleration, assuming an optimized hardware design that capitalizes on the inherent parallelism of FPGAs.
FPGA Type / Design Flow | Estimated Latency (per trace) | Required Buffer Size (samples) |
Low-end FPGA (Zynq) | 11.17 s | 2,792 samples |
Mid-range FPGA (Arria 10) | 4.78 s | 1,195 samples |
High-end FPGA (Alveo U280) | 2.23 s | 558 samples |
High-end (INT8-Optimized) | 1.12 s | 280 samples |
Here, the row labeled High-end (INT8-Optimized) represents a design flow that leverages 8-bit quantization on a device similar to other high-end FPGAs, rather than indicating a distinct family of boards. Narrower data paths in an INT8 pipeline can further reduce latency and resource utilization, at the expense of additional quantization efforts.
Even a low-end FPGA such as a Zynq device could reduce the baseline GPU latency from down to approximately , significantly lowering the buffering requirement. However, whether this meets “real-time constraints” depends on the total system requirements, including any overhead from I/O, external memory transfers, and other pipeline stages. In large-scale or power-sensitive deployments, factors such as cost, power consumption, and design complexity also guide the choice of FPGA platform and quantization scheme. The final selection must balance these considerations to ensure both feasibility and efficiency for the target application.
To fully enable real-time classification of cosmic-ray radio pulses, a dedicated FPGA-based framework must balance latency, resource utilization, and quantization precision while integrating seamlessly with experimental data streams. The results in Table 1 demonstrate that even low-end FPGA platforms can achieve substantial latency reductions, but practical deployment also requires optimizing DSP and memory utilization to handle high-rate data streams efficiently. Implementing 8-bit quantization (INT8) can further reduce both inference latency and resource consumption, although additional effort is needed to mitigate potential accuracy loss. A real-time classification system must also account for I/O bandwidth constraints, buffering mechanisms, and event prioritization to avoid pipeline stalls. Future work will focus on refining this framework by incorporating real-time data acquisition, firmware-level optimizations, and adaptive triggering strategies, ensuring a robust and scalable AI-driven self-triggering system for next-generation cosmic-ray experiments.
Future work will extend beyond optimizing the model for FPGA deployment while preserving classification performance. A key focus will be on benchmarking against alternative AI architectures, refining network optimization techniques, and expanding the training dataset to enhance generalization. Incorporating extensive air shower (EAS) simulations with state-of-the-art tools such as CoREAS will provide a more comprehensive and physically motivated dataset. Additionally, integrating experimental data from particle detector-triggered events will create an unprecedentedly realistic training set. Our preliminary results indicate that training on high-amplitude datasets enables the detection of even the faintest radio pulses—signals that have remained undetected by traditional methods to date. By leveraging these advancements, we aim to push the boundaries of real-time cosmic ray detection, making AI-driven self-triggering a transformative technology for next-generation astroparticle physics experiments.
6 Conclusion
In this work, we have demonstrated that artificial intelligence can dramatically enhance the self-triggering capability of radio detectors for extensive air showers (EAS). By training on large volumes of measured noise data and injected cosmic-ray-like pulses, our deep learning pipeline—incorporating both an autoencoder-based denoiser and a fully convolutional classifier—achieves excellent separation of genuine signals from background. Over a wide range of signal-to-noise ratios, the approach retains efficiencies in excess of 99% while simultaneously suppressing false positives down to sub-percent levels. Even in an urban-like noise environment, with RFI sources and transient interference, the model maintains robust performance, indicating strong potential for deployment under realistic conditions. Crucially, the flexibility of our inference stage enables an operating point where false positives can be entirely eliminated (as observed in our large-scale noise dataset), albeit at a modest reduction in true-positive rate to roughly 72%. This result opens the door to single-station cosmic-ray detection purely by radio triggers, removing the need to rely on any particle detectors for trigger generation. Although multiple-station coincidences remain a valuable method for additional false-positive suppression, these findings show that a single antenna unit could be made sufficiently selective to capture genuine air-shower events in high-noise environments. A key outcome of this study is the successful demonstration of latency estimates that make real-time triggering feasible. Our analysis of FPGA implementations shows that when using optimized designs and dataflow pipelines, inference times can reach the microsecond regime. This aligns well with the strict timing requirements of large-scale radio observatories, where trigger decisions must be made rapidly to record short-lived air shower signals. The ability to process data in-stream, directly on reconfigurable hardware, underscores the practical viability of these AI-driven methods for next-generation detectors. Overall, these results address one of the longstanding challenges in radio-based cosmic ray detection: reducing or eliminating the need for external triggers from particle or fluorescence detectors. An autonomous trigger operating purely on radio data lowers complexity and costs, while also unlocking sensitivity to high zenith angles and potentially exotic UHE particle interactions. Moving forward, further integration with simulation tools (e.g. CoREAS) and additional real shower data will refine the models’ accuracy and increase robustness to a wider spectrum of signal morphologies. The high efficiency and low false trigger rates shown here pave the way for a new class of AI-based, fully self-triggered radio arrays, promising both improved coverage and enhanced scientific reach in ultra-high-energy cosmic ray research.
7 Discussion
The results presented in this study highlight several key insights into the performance and robustness of our AI-driven self-triggering framework for radio detectors. Beyond achieving high detection efficiencies and minimal false-positive rates, a deeper examination of the model’s behaviour reveals its ability to generalise well across different noise conditions, preserve signal fidelity through denoising, and maintain robustness in realistic deployment scenarios.
A notable finding is the classifier’s ability to maintain high purity on non-denoised background traces, even though it was never trained on them. Because the classifier was trained exclusively on denoised signal and background data, one might have expected it to misinterpret raw background as containing signals. However, testing showed that it continued to distinguish signal from background with high accuracy, suggesting it was not misled by potential artifacts introduced by the denoiser on background samples. This indicates that the combined denoiser–classifier pipeline does not merely adapt to denoiser-induced modifications but instead learns robust, physically meaningful pulse features, demonstrating strong generalization across different noise conditions. Furthermore, these results suggest that mixing raw and denoised data in the training set may be an intriguing avenue for future exploration, as it could further improve robustness to a wider range of conditions.
We also investigated the effect of narrow-band RFI suppression. For the datasets discussed in this work, we did not initially remove strong RFI components. However, we produced an alternate noise dataset in which these narrow-band RFI components were suppressed prior to pulse injection. Because suppressing RFI lowers the noise floor by about 17 dB, we scaled down the injected pulses by a matching factor to preserve the same relative signal-to-noise ratio (SNR) distribution that our model was originally trained on. As expected, the classifier’s performance remained virtually unchanged, since the effective classification task difficulty was not altered. Nevertheless, eliminating RFI could simplify the denoiser and classifier in a real-world deployment, as it reduces background complexity. A simpler network architecture may lead to faster inference times and lower computational costs, paving the way for real-time triggering with reduced latency.
Another key consideration for real-world applications is the resilience of our pipeline to variations in noise characteristics. Our tests in an urban-like interference environment suggest that the denoiser–classifier chain is learning fundamental signal-versus-noise traits rather than overfitting to specific noise profiles. Further studies could systematically probe this robustness using augmented training datasets (e.g., artificially injecting varied noise profiles) or by testing on independent, geographically diverse datasets.
Beyond classification accuracy, latency remains critical for real-time self-triggering applications. Preliminary estimates indicate that microsecond inference is feasible on FPGA platforms, but hardware-specific optimisations are needed to confirm these figures in practice. Refining the neural network architectures to strike an optimal balance between complexity and speed could ensure that real-time decision-making constraints are met without compromising accuracy.
Looking ahead, our framework could be extended by incorporating CoREAS-simulated pulses covering a wide range of air shower scenarios, including variations in energy, zenith angle, and primary composition. Training on these simulated signals would allow the model to learn a more comprehensive representation of shower-induced radio emission, likely improving its ability to detect real cosmic-ray pulses in challenging regimes. This approach would also provide a clear path for systematically evaluating the pipeline’s performance under diverse air shower conditions.
Finally, although our results demonstrate the feasibility of single-antenna triggering, integrating AI-driven self-triggering with multi-station coincidence algorithms is an exciting possibility. Cross-station coincidence can lower false-positive rates while preserving sensitivity to rare events, including highly inclined cosmic rays that might elude traditional triggers.
In summary, our findings underscore the transformative potential of deep learning for self-triggered radio detection of extensive air showers. By preserving robust generalisation, demonstrating strong performance even with various noise conditions, and operating at latencies amenable to real-time applications, AI-based triggering stands poised to play an integral role in next-generation cosmic-ray observatories. Future work will focus on refining model architectures, evaluating performance across different noise environments, integrating large-scale radio arrays, and exploring how single-station AI triggers may combine with multi-station coincidence for enhanced detection confidence.
Acknowledgments
The author gratefully acknowledges the support of the Electronics Laboratory (E-Lab) of the Physics Department at Universität Siegen for their assistance in designing and producing the electronics used in this research. Their contributions facilitated the development and testing of the hardware components, streamlining the experimental setup.
This research made use of open-source software, including PyTorch and various scientific computing libraries in Python, such as NumPy, SciPy, and Matplotlib. The author acknowledges their developers for providing these valuable tools, which played a crucial role in data processing, model training, and analysis.

### v3

@@
 Abstract
-Cosmic-ray detection with radio antennas has traditionally depended on external triggers from particle detectors, constraining sensitivity and increasing complexity.
-Previous attempts at fully standalone, radio-only triggers have often failed under intense radio frequency interference, making genuine air-shower signals difficult to isolate.
-We present a proof-of-principle artificial intelligence-based self-triggering system that overcomes these limitations.
-By training a deep learning model on both real noise data and injected cosmic-ray-like pulses, we achieve an exceptionally low false-positive rate alongside high detection efficiency.
-Configurable operating points can suppress false positives below 0.01% while retaining more than 88% of genuine signals, and can even eliminate false positives entirely at a modest reduction in signal efficiency.
-This flexibility makes single-station cosmic-ray detection feasible without requiring external trigger inputs.
-Applying our approach to real-world noise conditions reduces the initial false-positive event rate by several orders of magnitude, supporting large-scale deployments.
-Extrapolation to dedicated hardware implementations, such as FPGAs, indicates that sub- inference times are achievable, enabling real-time autonomous triggering.
-These results highlight the transformative potential of artificial intelligence for enhancing radio detection sensitivity and inaugurate a new generation of fully self-triggered cosmic-ray observatories.
+Autonomous self-triggering for radio detection of extensive air showers remains a long-standing challenge, particularly in environments dominated by strong and variable radio-frequency interference.
+Current radio arrays usually rely on external particle-detector triggers: while this lowers thresholds for vertical showers, it excludes very inclined events, where radio detection is uniquely powerful because the particle cascade is absorbed while the radio pulse remains measurable.
+In this work, I present a proof-of-principle study showing that deep-learning-based triggering can overcome these limitations, operating robustly under realistic high-interference conditions and within the strict latency constraints of large-scale observatories.
+Using measured noise traces combined with simulated cosmic-ray pulses, a fully convolutional network was trained and optimised for MHz-scale trigger-trial rates, and its performance was compared to a simple threshold-based trigger, which proved markedly less efficient at the same false-positive rate.
+The trained model was then quantised with HLS4ML and synthesised with Vitis HLS for multiple FPGA targets.
+The quantised implementation preserves the floating-point model’s physics performance (, ) while achieving microsecond-scale inference latencies.
+These results show that modern FPGA-based AI inference can provide low-latency, radio-only triggering for next-generation cosmic-ray observatories, particularly in regions with strong and variable radio backgrounds.
+This capability unlocks access to weak and highly inclined air-shower signals—even in noisy environments—thereby broadening the energy range, extending sky coverage, and opening the path toward ultra-high-energy neutrino detection with ground-based radio arrays.
 1 Introduction
@@
 As a result, these efforts faced significant challenges and did not produce established scientific results, underscoring the need for more advanced signal-processing strategies.
+Notably, the ANITA [10] and TREND [11, 12] experiments successfully implemented autonomous radio self-triggering.
+ANITA, a balloon-borne radio interferometer, operated in the stratosphere, where the continuous background was comparatively quiet but impulsive anthropogenic RFI events were still present and were identified and rejected using polarisation information in subsequent analysis.
+TREND, by contrast, deployed a ground-based array in a remote, radio-quiet region of China, where ambient noise levels were exceptionally low.
+In both cases, comparatively favourable noise conditions were crucial in enabling effective self-triggering with simple threshold-based approaches—conditions that are significantly more challenging in the high-interference environments considered in this work.
 Recent developments in modelling EAS radio emissions, combined with experimental and simulation studies, have provided a deeper understanding of characteristic pulse shapes and ground radiation patterns [2].
@@
 Recent studies show that deep learning models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), can effectively distinguish cosmic-ray signals from background noise, significantly reducing detection thresholds.
-These techniques have been successfully applied to identify air-shower radio pulses in offline analyses [10], demonstrating their potential to enhance signal selection.
+These techniques have been successfully applied to identify air-shower radio pulses in offline analyses [13, 14, 15], demonstrating their potential to enhance signal selection.
 Recent efforts, such as those in the GRAND experiment, continue to investigate self-triggering using AI-based methods, including CNNs.
@@
 These advancements highlight the potential of FPGA-based AI techniques in handling complex, high-rate signal environments, which are also crucial for enhancing the efficiency and accuracy of cosmic-ray self-triggering.
-In this work, we investigate the feasibility of FPGA-based AI self-triggering for radio cosmic-ray detection.
-We begin by training and evaluating our deep neural network on Metal Performance Shaders (MPS) with data that includes experimentally measured background noise, then estimate their potential performance on FPGAs.
-While high-level synthesis (HLS) frameworks such as HLS4ML [14] and Xilinx Vitis AI [15] offer streamlined approaches to deploying AI models on FPGAs, we consider their integration a logical extension of this proof-of-concept study rather than part of our immediate scope.
-Our method comprises a fully convolutional autoencoder that denoises radio traces by suppressing background noise while retaining cosmic-ray pulses, followed by a fully convolutional classifier (FCN) that differentiates cosmic-ray signals from noise.
-By leveraging real data with measured background noise, we enhance the robustness of our approach in practical deployments.
-We also address critical issues associated with FPGA-based AI, including quantization strategies, resource utilization, and latency reduction, and propose a framework for real-time classification of cosmic-ray radio pulses.
-This proof-of-principle demonstration paves the way for future work incorporating more advanced HLS techniques to further optimize deployment on FPGA hardware.
+In this work, I investigate the feasibility of FPGA-based AI self-triggering for radio detection of EAS.
+I begin by training and evaluating a fully convolutional network on GPU using a dataset that combines experimentally measured background noise with simulated cosmic-ray pulses generated by the MGMR3D framework [19, 20, 21, 22].
+The trained model is then quantised with hls4ml [23, 24, 25] to fixed-point precision and synthesised using Vitis HLS 2023.2 [26] for multiple FPGA targets, allowing direct assessment of post-synthesis latency and resource utilisation.
+By leveraging measured noise from a high-interference environment, the approach is designed for robustness in realistic deployment conditions.
+I address key considerations for FPGA-based AI triggers, including quantisation, latency optimisation, and DSP-constrained architectures, and demonstrate that sub-microsecond inference latencies are achievable within realistic resource budgets.
 The implementation of AI-accelerated FPGA-based self-triggering represents a significant advancement in radio detection technology.
-By reducing reliance on external detector triggers and increasing detection efficiency, this approach enables greater autonomy for radio observatories, facilitating the detection of rare and high-energy events, including ultra-high-energy neutrinos.
-Our results demonstrate the feasibility of deploying AI-driven self-triggering at the detector level, marking an important step toward fully autonomous, real-time radio detection systems for astroparticle physics.
+By reducing reliance on external detector triggers and enabling high-efficiency, low-latency signal identification, this work opens the path toward autonomous operation of next-generation radio observatories.
+It marks the first demonstration, in the context of radio-detector self-triggering, of a complete workflow from model training to FPGA-synthesised deployment using hls4ml, providing a solid foundation for future developments in real-time cosmic-ray radio detection.
 2 Experimental Setup
@@
 Given the high variability of the urban noise environment, the dataset presents a challenging test case for AI-driven signal detection methods.
-This setting introduces additional challenges, including increased transient interference and strong radio frequency interference (RFI).
+This setting introduces additional challenges, including increased transient interference and strong RFI.
 Two prominent narrowband RFI sources were identified at approximately and , along with other narrowband and transient RFI.
@@
 The results presented in the following section reflect model performance when both training and testing were conducted on data sets containing these RFIs.
-Additionally, we examine how performance changes when these narrowband RFIs are mitigated prior to training and validation, providing insights into the impact of RFI suppression on the learning process, which we will report in the discussion section 7.
-To evaluate the AI model’s performance, signal-like pulses were injected into measured noise traces at random positions, covering a broad range of amplitudes.
-The pulse power was scaled relative to the noise power, which was calculated using a pre-filtered version of each trace.
-This ensured that the injected pulses maintained a well-controlled signal-to-noise ratio (SNR) within a user-defined range while accurately reflecting real-world noise conditions.
-However, to prevent double filtering of the noise, which could distort its characteristics, the pulses were injected into the original (unfiltered) traces.
-This step also eliminated systematic biases by ensuring that the statistical properties of the background noise remained unaffected by the injection process.
-To maintain consistency in data processing, the digital bandpass filter was applied only after pulse injection, ensuring that both noise-only and signal-injected datasets underwent identical filtering and conditioning.
-This approach was chosen over COREAS simulations [16] to establish a proof of principle based purely on signal strength, avoiding dependencies on the absolute energy scale, antenna gain pattern, and shower inclination, which introduce additional uncertainties.
-By injecting short pulses into measured noise traces from a real-world, high-interference environment, we enable a focused evaluation of the AI model’s ability to detect transient signal shapes under conditions where traditional threshold-based triggering systems struggle or fail entirely due to overwhelming radio frequency interference (RFI).
-This approach enables a systematic assessment of the model sensitivity in varying SNRs while ensuring robustness under challenging conditions.
-By directly testing and refining AI-based techniques on measured noise traces—without relying on an external event-reconstruction framework or requiring a particle detector—it provides a controlled and scalable test case.
-Although it does not fully replicate all details of cosmic-ray pulse shapes, much of which may already be washed out by bandpass filtering, it offers a robust estimate of the false positive rate of AI-based triggering methods in a realistic noise environment.
+To evaluate the AI model’s performance under realistic and diverse signal conditions, I generated air-shower radio pulses using the MGMR3D simulation framework [19, 20, 21, 22], which reproduces the time-domain waveform and frequency content expected from extensive air showers for a wide range of shower geometries and primary energies.
+The simulated dataset spans zenith angles from to in steps, uniformly distributed random azimuths, and primary energies between and .
+For each geometry, was varied randomly between 600 and to cover the natural spread observed in air-shower development.
+Signals were simulated for antenna positions at distances of to from the shower core, ensuring realistic variation in pulse amplitude, shape, polarity, and spectral content across the dataset.
+In total, this resulted in the generation of over 78,000 simulated radio pulses.
+These pulses, already encompassing a wide range of shapes and amplitudes from the simulated variation in energy, geometry, values and antenna–core distance, were injected at random times into measured noise traces from a high-interference environment.
+To control the signal-to-noise ratio (SNR) within a user-defined range—without performing an absolute amplitude conversion from simulated electric-field strength to the measured detector response—I rescaled each simulated pulse relative to the noise power, calculated from a pre-filtered version of the corresponding trace.
+To preserve the statistical properties of the noise and avoid any distortion from repeated filtering, pulses were injected into the original (unfiltered) traces.
+The same digital bandpass filter was then applied to both noise-only and signal-injected traces, ensuring identical processing and conditioning across the dataset.
+By combining realistic MGMR3D-generated air-shower pulses with measured noise, this approach enables a robust and representative evaluation of AI-based triggering performance.
+It captures the natural diversity of pulse shapes expected in real events while providing controlled SNR conditions.
+This ensures that the results are directly relevant for estimating the achievable false-positive rate under realistic and challenging RFI conditions.
 2.1 Training and Validation Datasets
@@
 For the initial analysis, only a bandpass filter was applied, with no additional filtering.
-However, we later incorporate further filtering to suppress prominent RFI sources in the 30– range for comparison, the results of which are discussed in section 7.
 Training and Cross-Validation Datasets: The dataset used for training and model tuning was randomly split into two equal parts:
 Training Dataset (50%): Comprised 15,000 time series traces, each containing 128 samples recorded at a sampling rate of , consistent with standard practices in radio-based air shower detection.
-It included an equal distribution of 5,000 pure noise traces (background), 5,000 traces of isolated EAS-like pulses (pure signal), and 5,000 traces with EAS-like pulses injected into noise (signal).
+It included an equal distribution of 5,000 pure noise traces (background), 5,000 traces of simulated EAS pulses (pure signal), and 5,000 traces with simulated EAS pulses injected into noise (signal).
 This setup ensured that the model effectively learned to distinguish genuine signals from background noise.
@@
 These independent validation data sets were crucial to assessing the model’s ability to generalize to previously unseen data and to reliably detect signals under varying noise conditions, particularly at low SNR.
+To quantify the SNR in a way that is robust to real detector conditions, I adopt a trace-wise definition using pure signal and pure noise traces (background) as inputs.
+Specifically, I define SNR as the squared peak amplitude of the Hilbert envelope of a pure signal trace, divided by the variance of a matched pure noise trace.
+This avoids ambiguity over whether the "signal" contains background or not.
+Formally, I compute
+where is the maximum amplitude of the analytic signal (obtained via the Hilbert transform) and is the noise power computed as the variance of a pure noise trace.
+This approach emphasises the prominence of the signal over the background without being sensitive to specific waveform shapes.
+To provide a baseline for background-only traces, I apply the same method to pure noise traces, treating them as if they contained a signal.
+Specifically, I compute the Hilbert envelope of each noise trace and take the maximum value as a hypothetical signal peak, then divide by the noise variance.
+This procedure captures spurious fluctuations in the noise that might mimic signal-like features.
+As a result, I obtain a distribution of apparent SNR values for background, which are typically lower than true signals but may overlap in some cases.
+This overlap justifies the use of logarithmic scaling and motivates the application of probabilistic, rather than deterministic, classification thresholds.
 Figure 2 illustrates the SNR distributions for background and signal traces across Validation Dataset 1, which closely follows the distribution of the training dataset, and Validation Dataset 2, which consists of low-amplitude signals.
 This comparison provides insights into the differences in signal detectability across varying SNR conditions.
+In the SNR histograms, I report the base-10 logarithm of the computed SNR.
 3 Methodology
-We implemented a deep learning-based pipeline for denoising and classifying radio time series in the context of self-triggering for EAS events. The pipeline consists of two primary components:
-1.
-A CNN-based denoiser to suppress noise in the input signals.
-2.
-A fully convolutional classifier (FCN) to distinguish between signal and background in a binary classification task.
-These models operate sequentially, where the denoiser first refines raw data before passing it to the classifier for event discrimination.
+The goal of my approach is to design, train, and evaluate a neural network classifier capable of robustly detecting air-shower radio pulses in realistic, high-interference noise conditions, while ensuring compatibility with efficient FPGA deployment.
+The methodology comprises two main components: (1) the design of a fully convolutional network architecture optimised for both classification performance and hardware constraints, and (2) the training and optimisation procedure used to achieve high detection efficiency with minimal false-positive rate.
 3.1 Model Architecture
-3.1.1 CNN-Based Denoiser
-The denoiser is a fully convolutional autoencoder designed to suppress noise and enhance cosmic ray pulses in noisy radio time series..
-It follows an encoder-bottleneck-decoder structure with a skip connection to retain fine-grained features.
-Encoder: Two convolutional layers extract multi-scale features, followed by a single max pooling layer that reduces resolution by a factor of two.
-Bottleneck: A deeper convolutional block refines feature representations at the reduced resolution.
-Decoder: An upsampling layer restores the original resolution, followed by convolutional layers that refine the denoised signal.
-A skip connection concatenates high-resolution features from the encoder to enhance detail recovery.
-The denoiser learns from a dataset where the inputs are noisy traces containing EAS-like pulses (signal), and the targets are the corresponding pure EAS-like pulses (pure signal).
-The training process minimises the Mean Squared Error (MSE) loss between the clean target and the denoised prediction :
-This architecture ensures that the model preserves cosmic ray features while effectively filtering out background noise.
-3.1.2 Fully Convolutional Classifier (FCN)
-The FCN classifier classifies time series into background or signal, where signal refers to a noisy trace containing an EAS pulse. It includes:
-Three convolutional blocks with batch normalization, dropout, and max pooling.
-Global average pooling (GAP) to reduce spatial dimensions while retaining features.
-A fully connected layer with a sigmoid activation function for binary classification.
-The classifier is trained using the Binary Cross-Entropy (BCE) loss:
-where is the true label and is the predicted probability.
-3.2 Training and Optimization
-Both models were trained using the Adam optimizer, with task-specific loss functions, following a two-step training approach:
-1.
-Denoiser Training:
-Trained on traces with pulses injected into noise (signal) and pure signal.
-Optimised to minimize Mean Squared Error (MSE) loss, enhancing signal quality while suppressing noise.
-2.
-Classifier Training:
-Trained on denoised signal traces (output of the denoiser) alongside denoised background traces.
-Optimized to minimize Binary Cross-Entropy (BCE) loss, ensuring accurate discrimination between signal-containing and noise-only traces.
-The model was trained for 15 epochs with a batch size of 64 and a learning rate of 0.001, using the Training Dataset.
-Its generalization performance was monitored with the Cross-Validation Dataset described in Section 2.1.
-The model was implemented using PyTorch with support for MPS acceleration.
-Training was conducted using the Adam optimizer, and model performance was evaluated on the cross-validation dataset using classification accuracy, precision, recall, and F1-score.
-4 Model Performance Analysis
-The first two plots in Figure 3 (top row) illustrate the probability distributions of sigmoid prediction for validation datasets 1 and 2, which differ in their SNR distributions.
-In Validation Dataset 1, where the SNR distribution closely resembles that of the training set (see Section 2.1), the model exhibits a strong separation between background and signal traces.
-Most noise traces receive lower probabilities, while signal-containing traces are predominantly assigned higher values.
-In contrast, Validation Dataset 2 contains traces with significantly lower SNRs, commonly below the noise floor.
-While the model still distinguishes between the two classes, the classification performance degrades: the distribution of signal-containing traces is more spread out, with a noticeable tail extending toward lower probabilities.
-This shift indicates a reduction in signal efficiency, as weaker signals are increasingly misclassified due to their low SNR.
-Beyond assessing performance across different SNRs, it is also essential to evaluate the model’s behaviour in environments dominated by overwhelming background noise, where false positives become a critical concern.
-To investigate this, we introduce Validation Dataset 3, which consists of one million background traces.
-The bottom row of Figure 3 explores the model’s response to this dataset by comparing background traces from Validation Dataset 3 with signal traces from Validation Dataset 1.
-The sigmoid probability distribution (bottom left) shows that while most noise traces receive low probabilities, the large dataset size results in a fraction extending into the signal region.
-This highlights the need to quantify the false trigger rate under realistic high-background conditions.
-The final plot (bottom right) presents the distributions of the model’s logits, i.e.
-the raw scores before applying a probability function.
-These distributions reveal a clear separation between background traces from Validation Dataset 3 and signal traces from Validation Dataset 1.
-Using logits instead of sigmoid probabilities significantly improves false positive rejection, albeit at the cost of true positive efficiency.
-This trade-off is particularly relevant when designing self-triggering strategies in noisy environments.
-4.1 Performance Metrics and Trigger Rates
-Following the qualitative assessment in Figure 3, we systematically evaluate the model’s performance across different validation datasets.
-This analysis focuses on three key aspects: trigger efficiency, robustness to low-SNR signals, and false trigger rate in high-background conditions.
-The results are presented as follows.
-4.1.1 Trigger Efficiency in Controlled Conditions (Validation Dataset 1)
-True Positives: 99.93%
-False Positives: 0.07%
-A simple probability threshold is sufficient for efficient signal selection in this case.
-4.1.2 Robustness to Weaker Pulses (Validation Dataset 2)
-True Positives: 87.96%
-False Positives: 0.08%
-Despite the increased challenge, the model effectively distinguishes signal-containing traces, demonstrating its robustness to weaker pulses.
-The reduction in efficiency is expected as most of these signals fall outside the training range.
-Although further tuning could enhance performance, the primary objective here is to assess the model’s ability to generalise rather than optimise for specific low-SNR cases.
-4.1.3 False Trigger Rate Estimation in a High-Noise Environment (Validation Dataset 3)
-To assess the model’s performance in the presence of overwhelming background noise, we evaluated it on Validation Dataset 3.
-This dataset simulates realistic high-RFI conditions, such as those found in urban environments (e.g., LOFAR, LOPES) or certain AERA stations [7].
-The primary objective of this evaluation is to quantify false trigger rates in such environments. The results are as follows:
-True Positives (Initial Threshold): 99.93%
-False Positives: 768 per 1M events (0.08%), which translates to approximately 800 Hz first-level trigger (FLT) rate in a 1 MHz trigger scenario.
-Applying confidence-based selection criteria (referred here as entropy cut) significantly reduces the false positive rate while maintaining high true positive efficiency.
-Entropy Cut: 99.35% True Positives, 79 False Positives per 1M events (0.008%), corresponding to approximately 80 Hz false triggers in a 1 MHz trigger scenario.
-The FLT rate is already suitable for a station-level trigger and provides a strong foundation for further refinement.
-A subsequent second-level trigger (SLT) stage, utilizing array-level coincidence searches, could further suppress false positives while preserving efficiency.
-To explore the feasibility of detecting cosmic-ray events at a single station, we further applied logits-based cuts, which resulted in:
-Logit Cut: 72.4% True Positives, 0 False Positives, completely eliminating false triggers in this dataset.
-This analysis shows that entropy-based selection effectively reduces the trigger rate to an acceptable level for an FLT, while still maintaining high detection efficiency.
-Further refinement at the SLT stage, through array-level coincidence searches, could optimise false positive rejection even further.
-On the other hand, applying logits-based cuts enables the possibility of detecting cosmic-ray events at a single station, drastically suppressing false positives, with none observed in our dataset.
-While this comes with some reduction in signal efficiency, it remains within a reasonable range, making single-station detection feasible in certain scenarios.
-4.1.4 Example of a Low-Amplitude Signal Classification
+The classification network is a fully convolutional 1-D architecture for radio traces with one input channel.
+It contains approximately trainable parameters and was designed to maintain classification performance comparable to earlier floating-point models while ensuring that each convolutional or linear layer contains fewer than weights—a practical constraint arising from Vitis HLS synthesis failures when unrolling larger layers due to internal 12-bit address limitations in the memory architecture of Vitis HLS, which restrict access to weight arrays exceeding 4096 entries.
+This constraint facilitates efficient FPGA implementation in later stages of the work without excessive on-chip memory use or pipeline latency.
+The network consists of three convolutional blocks followed by global average pooling and a fully connected output layer.
+Batch normalisation (BN) is applied after each convolution, ReLU activations follow all but the final convolution in each block, and dropout is applied after each pooling stage.
+The first block applies a 1D convolution with kernel size 3, 1 input channel, and 16 output channels (),
+where denotes the ReLU activation. This is followed by max-pooling with factor 2 and dropout with rate 0.3.
+The second block begins with a 1D convolution of kernel size 3 that increases the number of channels from 16 to 32, followed by batch normalisation and ReLU activation:
+This is followed by a residual sub-network composed of lightweight 1D convolutions:
+where consists of two 1D convolutional layers with reduced intermediate channel width, each followed by batch normalisation and activation.
+This design reduces the number of parameters while preserving the benefits of residual learning.
+The block ends with max-pooling (factor 2) and dropout with rate 0.3.
+The third block increases the number of channels from 32 to 64 using a bottleneck-style expansion module composed of three 1D convolutions:
+where and are 1D convolutions with kernel size 1, and is a 1D convolution with kernel size 3.
+Let denote the batch normalisation applied after the convolution.
+The first two convolutions are each followed by batch normalisation and ReLU activation; the final convolution is followed only by batch normalisation.
+This design enables parameter-efficient widening by reducing the number of channels in the intermediate layers, avoiding the cost of a wide convolution.
+The block concludes with max-pooling (factor 2) and dropout with rate 0.3.
+The classification head applies global average pooling across the temporal dimension, reducing the output tensor to a 64-dimensional feature vector by averaging across the temporal axis.
+A fully connected layer (641) produces the binary classification logit without an activation function, as the model is trained using a loss function that incorporates the sigmoid non-linearity.
+This description provides all hyperparameters, activation placements, and custom-module definitions required to reproduce the model exactly.
+3.2 Training and Optimisation
+The previously described architecture was trained to classify traces as either pure noise (background) or containing simulated cosmic-ray pulses (signal), as defined in Section 2.1.
+The loss function was the binary cross-entropy with logits,
+where is the model’s raw output (logit), the binary target label, the sigmoid function, and the mini-batch size.
+The model was optimised using the Adam algorithm [27] as implemented in PyTorch, with a fixed learning rate of , and default values for the exponential decay rates and numerical stability parameter: , , and .
+Training was performed for 15 epochs, and the model state corresponding to the lowest validation loss was retained for evaluation.
+The dataset used for model training and evaluation consisted of normalised, fixed-length radio traces drawn from two equally sized classes, corresponding to the training and cross-validation datasets defined in Section 2.1.
+The dataset was randomly shuffled and split into training and validation subsets.
+Training mini-batches were drawn from both classes in equal proportion to avoid class imbalance.
+To improve robustness for hardware-friendly deployment, unstructured magnitude pruning was applied to all convolutional layers.
+A polynomial schedule gradually increased weight sparsity from zero to 50% between early and mid-training epochs:
+where denotes the sparsity at epoch , is the target sparsity, and and are the pruning schedule boundaries.
+During training, pruning masks were dynamically updated, and after training completed, masks were removed to yield the final pruned model.
+After each epoch, the network was evaluated on the validation set without gradient updates.
+The primary selection criterion for saving the best model was the validation loss.
+For monitoring purposes during development, the area under the receiver operating characteristic (ROC) curve (AUC) was also computed using a fixed classification threshold of 0.5.
+The complete training pipeline, including dataset preparation, augmentation through pulse injection, pruning schedule, and validation, was implemented in PyTorch and executed on CUDA-enabled GPUs.
+4 Model Performance Analysis (Float Model)
+To evaluate the classifier’s final performance in a realistic triggering context, I consider signal traces from Validation Dataset 1 and background traces from Validation Dataset 3.
+Validation Dataset 1 contains noise traces with injected MGMR3D-simulated air-shower pulses at controlled SNRs matching the training distribution.
+Validation Dataset 3 consists of one million measured background traces recorded in a high-interference environment.
+The large sample size enables the evaluation of the model’s false-positive rejection efficiency in a highly challenging regime, where even a small false-positive rate would correspond to rates approaching the MHz range in a real-time self-trigger application.
+The main performance indicator is the receiver operating characteristic (ROC) curve (Figure 3), obtained from classifier outputs for signal traces from Validation 1 and background traces from Validation 3.
+The solid curve shows the trade-off between the true positive rate (signal efficiency) and the false positive rate as the classification threshold is varied, with an area under the curve (AUC) of 0.969.
+The vertical red dashed line marks the chosen operating point at a false positive rate of , illustrating that the model maintains high efficiency while suppressing false triggers to extremely low rates.
+While the ROC analysis focuses on the nominal SNR distribution, model performance was also evaluated on Validation Dataset 2, where most injected pulses have SNRs well below the noise floor.
+At a fixed false positive rate of , the signal efficiency decreases from on Validation 1 to on Validation 2.
+This drop reflects the expected trend toward lower classification scores for weaker signals.
+These results provide a first indication that the network retains some discriminative power in the low-SNR regime.
+4.1 False Trigger Rate Estimation in a High-Noise Environment
+To assess the model’s robustness in realistic high-noise conditions, I evaluated it on Validation Dataset 3.
+This dataset reflects high-RFI environments such as those encountered in urban settings (e.g., LOFAR, LOPES) or at certain AERA stations [7].
+Such conditions are representative of operational scenarios in which a self-triggering station must continuously process high-rate background waveforms, potentially at sampling rates of up to trigger trials.
+In this regime, even a small false positive rate (FPR) translates into a substantial trigger load: for example, an corresponds to first-level trigger (FLT) rate per station.
+From the ROC analysis (Figure 3), the model achieves a signal efficiency of at this target operating point.
+The FLT rate is already suitable for a station-level trigger in a cosmic-ray radio array and provides a strong foundation for further refinement.
+In practice, a subsequent second-level trigger (SLT) stage—based on array-level coincidence requirements—can further suppress false positives while preserving nearly all genuine events.
+Such multi-stage triggering strategies are widely used in cosmic-ray experiments (e.g., Auger, AERA, GRAND) and are essential for handling high raw data rates in a scalable and efficient manner.
+For applications where further reduction of the FLT rate is desired, the threshold can be tightened to , corresponding to roughly at the station level.
+This operating point still retains a signal efficiency of and may be useful in scenarios where exceptionally low background rates are advantageous.
+In practice, already offers a good balance between efficiency and manageable trigger rates, with the lower-FPR setting available as an option when background suppression takes priority.
+4.2 Example of a Low-Amplitude Signal Classification
 Figure 4 illustrates a representative trace containing a faint signal pulse correctly classified by the model.
@@
 This result highlights the model’s ability to recover weak air shower signals that would otherwise remain undetectable using traditional threshold-based methods.
-5 Estimated Latency on FPGA Platforms
-To evaluate the feasibility of real-time deployment, the model’s inference latency was first measured on an Apple Silicon GPU using Metal Performance Shaders (MPS).
-Each inference took approximately per trace, which implies a required buffer of around 8,250 samples when sampling at .
-Since real-time streaming at this rate is continuous, reducing inference latency is essential to minimize buffering and avoid potential pipeline stalls.
-Moving from a GPU to an FPGA requires a different set of assumptions compared to traditional CPU-to-FPGA speed-up estimates.
-In particular, FPGAs allow for specialized parallelism and hardware-friendly dataflow optimizations that can significantly reduce latency if designed properly.
-Table 1 summarizes conservative latency estimates for different FPGA classes, along with the corresponding buffer sizes at .
-These values were derived by adjusting for GPU-to-FPGA acceleration, assuming an optimized hardware design that capitalizes on the inherent parallelism of FPGAs.
-FPGA Type / Design Flow | Estimated Latency (per trace) | Required Buffer Size (samples) |
-Low-end FPGA (Zynq) | 11.17 s | 2,792 samples |
-Mid-range FPGA (Arria 10) | 4.78 s | 1,195 samples |
-High-end FPGA (Alveo U280) | 2.23 s | 558 samples |
-High-end (INT8-Optimized) | 1.12 s | 280 samples |
-Here, the row labeled High-end (INT8-Optimized) represents a design flow that leverages 8-bit quantization on a device similar to other high-end FPGAs, rather than indicating a distinct family of boards.
-Narrower data paths in an INT8 pipeline can further reduce latency and resource utilization, at the expense of additional quantization efforts.
-Even a low-end FPGA such as a Zynq device could reduce the baseline GPU latency from down to approximately , significantly lowering the buffering requirement.
-However, whether this meets “real-time constraints” depends on the total system requirements, including any overhead from I/O, external memory transfers, and other pipeline stages.
-In large-scale or power-sensitive deployments, factors such as cost, power consumption, and design complexity also guide the choice of FPGA platform and quantization scheme.
-The final selection must balance these considerations to ensure both feasibility and efficiency for the target application.
-To fully enable real-time classification of cosmic-ray radio pulses, a dedicated FPGA-based framework must balance latency, resource utilization, and quantization precision while integrating seamlessly with experimental data streams.
-The results in Table 1 demonstrate that even low-end FPGA platforms can achieve substantial latency reductions, but practical deployment also requires optimizing DSP and memory utilization to handle high-rate data streams efficiently.
-Implementing 8-bit quantization (INT8) can further reduce both inference latency and resource consumption, although additional effort is needed to mitigate potential accuracy loss.
-A real-time classification system must also account for I/O bandwidth constraints, buffering mechanisms, and event prioritization to avoid pipeline stalls.
-Future work will focus on refining this framework by incorporating real-time data acquisition, firmware-level optimizations, and adaptive triggering strategies, ensuring a robust and scalable AI-driven self-triggering system for next-generation cosmic-ray experiments.
-Future work will extend beyond optimizing the model for FPGA deployment while preserving classification performance.
-A key focus will be on benchmarking against alternative AI architectures, refining network optimization techniques, and expanding the training dataset to enhance generalization.
-Incorporating extensive air shower (EAS) simulations with state-of-the-art tools such as CoREAS will provide a more comprehensive and physically motivated dataset.
-Additionally, integrating experimental data from particle detector-triggered events will create an unprecedentedly realistic training set.
-Our preliminary results indicate that training on high-amplitude datasets enables the detection of even the faintest radio pulses—signals that have remained undetected by traditional methods to date.
-By leveraging these advancements, we aim to push the boundaries of real-time cosmic ray detection, making AI-driven self-triggering a transformative technology for next-generation astroparticle physics experiments.
-6 Conclusion
-In this work, we have demonstrated that artificial intelligence can dramatically enhance the self-triggering capability of radio detectors for extensive air showers (EAS).
-By training on large volumes of measured noise data and injected cosmic-ray-like pulses, our deep learning pipeline—incorporating both an autoencoder-based denoiser and a fully convolutional classifier—achieves excellent separation of genuine signals from background.
-Over a wide range of signal-to-noise ratios, the approach retains efficiencies in excess of 99% while simultaneously suppressing false positives down to sub-percent levels.
-Even in an urban-like noise environment, with RFI sources and transient interference, the model maintains robust performance, indicating strong potential for deployment under realistic conditions.
-Crucially, the flexibility of our inference stage enables an operating point where false positives can be entirely eliminated (as observed in our large-scale noise dataset), albeit at a modest reduction in true-positive rate to roughly 72%.
-This result opens the door to single-station cosmic-ray detection purely by radio triggers, removing the need to rely on any particle detectors for trigger generation.
-Although multiple-station coincidences remain a valuable method for additional false-positive suppression, these findings show that a single antenna unit could be made sufficiently selective to capture genuine air-shower events in high-noise environments.
-A key outcome of this study is the successful demonstration of latency estimates that make real-time triggering feasible.
-Our analysis of FPGA implementations shows that when using optimized designs and dataflow pipelines, inference times can reach the microsecond regime.
-This aligns well with the strict timing requirements of large-scale radio observatories, where trigger decisions must be made rapidly to record short-lived air shower signals.
-The ability to process data in-stream, directly on reconfigurable hardware, underscores the practical viability of these AI-driven methods for next-generation detectors.
-Overall, these results address one of the longstanding challenges in radio-based cosmic ray detection: reducing or eliminating the need for external triggers from particle or fluorescence detectors.
-An autonomous trigger operating purely on radio data lowers complexity and costs, while also unlocking sensitivity to high zenith angles and potentially exotic UHE particle interactions.
-Moving forward, further integration with simulation tools (e.g.
-CoREAS) and additional real shower data will refine the models’ accuracy and increase robustness to a wider spectrum of signal morphologies.
-The high efficiency and low false trigger rates shown here pave the way for a new class of AI-based, fully self-triggered radio arrays, promising both improved coverage and enhanced scientific reach in ultra-high-energy cosmic ray research.
-7 Discussion
-The results presented in this study highlight several key insights into the performance and robustness of our AI-driven self-triggering framework for radio detectors.
-Beyond achieving high detection efficiencies and minimal false-positive rates, a deeper examination of the model’s behaviour reveals its ability to generalise well across different noise conditions, preserve signal fidelity through denoising, and maintain robustness in realistic deployment scenarios.
-A notable finding is the classifier’s ability to maintain high purity on non-denoised background traces, even though it was never trained on them.
-Because the classifier was trained exclusively on denoised signal and background data, one might have expected it to misinterpret raw background as containing signals.
-However, testing showed that it continued to distinguish signal from background with high accuracy, suggesting it was not misled by potential artifacts introduced by the denoiser on background samples.
-This indicates that the combined denoiser–classifier pipeline does not merely adapt to denoiser-induced modifications but instead learns robust, physically meaningful pulse features, demonstrating strong generalization across different noise conditions.
-Furthermore, these results suggest that mixing raw and denoised data in the training set may be an intriguing avenue for future exploration, as it could further improve robustness to a wider range of conditions.
-We also investigated the effect of narrow-band RFI suppression.
-For the datasets discussed in this work, we did not initially remove strong RFI components.
-However, we produced an alternate noise dataset in which these narrow-band RFI components were suppressed prior to pulse injection.
-Because suppressing RFI lowers the noise floor by about 17 dB, we scaled down the injected pulses by a matching factor to preserve the same relative signal-to-noise ratio (SNR) distribution that our model was originally trained on.
-As expected, the classifier’s performance remained virtually unchanged, since the effective classification task difficulty was not altered.
-Nevertheless, eliminating RFI could simplify the denoiser and classifier in a real-world deployment, as it reduces background complexity.
-A simpler network architecture may lead to faster inference times and lower computational costs, paving the way for real-time triggering with reduced latency.
-Another key consideration for real-world applications is the resilience of our pipeline to variations in noise characteristics.
-Our tests in an urban-like interference environment suggest that the denoiser–classifier chain is learning fundamental signal-versus-noise traits rather than overfitting to specific noise profiles.
-Further studies could systematically probe this robustness using augmented training datasets (e.g., artificially injecting varied noise profiles) or by testing on independent, geographically diverse datasets.
-Beyond classification accuracy, latency remains critical for real-time self-triggering applications.
-Preliminary estimates indicate that microsecond inference is feasible on FPGA platforms, but hardware-specific optimisations are needed to confirm these figures in practice.
-Refining the neural network architectures to strike an optimal balance between complexity and speed could ensure that real-time decision-making constraints are met without compromising accuracy.
-Looking ahead, our framework could be extended by incorporating CoREAS-simulated pulses covering a wide range of air shower scenarios, including variations in energy, zenith angle, and primary composition.
-Training on these simulated signals would allow the model to learn a more comprehensive representation of shower-induced radio emission, likely improving its ability to detect real cosmic-ray pulses in challenging regimes.
-This approach would also provide a clear path for systematically evaluating the pipeline’s performance under diverse air shower conditions.
-Finally, although our results demonstrate the feasibility of single-antenna triggering, integrating AI-driven self-triggering with multi-station coincidence algorithms is an exciting possibility.
-Cross-station coincidence can lower false-positive rates while preserving sensitivity to rare events, including highly inclined cosmic rays that might elude traditional triggers.
-In summary, our findings underscore the transformative potential of deep learning for self-triggered radio detection of extensive air showers.
-By preserving robust generalisation, demonstrating strong performance even with various noise conditions, and operating at latencies amenable to real-time applications, AI-based triggering stands poised to play an integral role in next-generation cosmic-ray observatories.
-Future work will focus on refining model architectures, evaluating performance across different noise environments, integrating large-scale radio arrays, and exploring how single-station AI triggers may combine with multi-station coincidence for enhanced detection confidence.
+4.3 Comparison with a Threshold-Based Trigger
+To provide a baseline against conventional self-triggering techniques, I implemented a simple SNR-threshold trigger.
+For each evaluation trace, I form the analytic signal via the Hilbert transform and define the score as in Eq.
+2.1.
+For background traces, the trace itself serves as this reference, so the largest envelope excursion is treated as a spurious candidate signal.
+The resulting distributions of scores for signal and background events allow the construction of a ROC curve by varying the threshold on the SNR score.
+To ensure a fair comparison, both the neural-network classifier and the SNR-threshold baseline were evaluated on the same unfiltered datasets, which contain the full interference environment as described in Section 2.1.
+In practice, threshold-based self-triggers are often preceded by frequency filtering to increase sensitivity, but such preprocessing would also be expected to benefit the neural network if applied consistently.
+Applying additional filtering only to the baseline would artificially simplify its input and alter the effective noise conditions relative to the classifier evaluation.
+For this reason, I restrict the comparison to identical unfiltered inputs for both methods.
+For a fixed input, ROC performance is invariant under any monotonic reparameterisation of the score.
+For example, using the peak-to-RMS ratio or the z-score is mathematically equivalent to our SNR definition up to a monotonic transform and therefore yields the same ROC curve within numerical precision.
+By contrast, using the raw peak voltage without noise normalisation performs significantly worse, as expected, because it does not account for fluctuations in the noise level.
+More elaborate variants (e.g.
+running-average baselines, short-time energy) remain threshold triggers in spirit but introduce method-specific hyper-parameters and design choices.
+To avoid cherry-picking, I therefore restrict the comparison to this minimal representative baseline.
+Figure 5 shows the ROC of the SNR-threshold trigger.
+The area under the curve is .
+At the operating point used throughout this work, corresponding to a false positive rate of (vertical red dashed line), the threshold-based trigger achieves a signal efficiency of only .
+For comparison, the neural-network classifier maintains an efficiency of under the same conditions (cf.
+Fig.
+3).
+This demonstrates a clear improvement of the AI-based approach over a traditional threshold trigger in the challenging high-interference environment considered here.
+4.4 Physics Implications of Improved Triggering
+Beyond the technical demonstration, it is instructive to consider the physics impact of lowering the trigger threshold in realistic, interference-rich environments.
+At the chosen operating point (false positive rate of ), the SNR-threshold baseline achieves a signal efficiency of only (Sec.
+4.3), whereas the neural-network classifier retains an efficiency of (Sec.
+4).
+This corresponds to an effective reduction of the detectable signal-to-noise threshold by more than a factor of two.
+Such an improvement has direct consequences for the reach of radio arrays.
+Lowering the trigger threshold extends the accessible energy range towards lower primary energies, where air-shower radio signals are intrinsically weaker.
+It also improves sensitivity to very inclined air showers, since the particle component is largely absorbed in the atmosphere while the radio emission remains strong.
+These events are particularly valuable for composition studies, for enlarging the sky coverage of ground-based observatories, and as a prerequisite for detecting ultra-high-energy neutrinos, which manifest as upward-going or near-horizontal showers.
+By enabling robust radio-only operation under high-background conditions, the method reduces reliance on external particle detectors, thereby simplifying array design and allowing sparser station layouts.
+This opens a path toward larger effective apertures at EeV energies, complementing and extending the capabilities of existing cosmic-ray and neutrino observatories.
+These considerations show that the proposed AI-based trigger is not only technically feasible on FPGA hardware, but also provides a tangible physics benefit by increasing the accessible parameter space of radio arrays in energy, inclination, and coverage.
+5 Quantisation and FPGA Implementation
+5.1 Quantisation with HLS4ML
+For real-time deployment, the trained floating-point classifier was converted to a fixed-point representation using the HLS4ML framework [23, 24, 25].
+Quantisation replaces floating-point multiplications and additions with fixed-point arithmetic, reducing resource usage and latency on FPGAs while maintaining compatibility with hardware synthesis workflows.
+The key challenge is to choose a precision that minimises resource cost without sacrificing classification performance.
+To determine a suitable fixed-point format for FPGA deployment, I first estimated the minimum required integer bit-width to avoid overflow.
+This was done by forwarding representative validation samples through the trained model while recording the maximum absolute activation values across all intermediate layers.
+Based on this analysis, I fixed the integer width at 5 bits.
+Subsequently, I scanned the total bit-width while keeping the integer part fixed, evaluating classification performance at each precision setting.
+Figure 6a shows the area under the ROC curve (AUC) as a function of total bit-width.
+Performance remains stable down to 13 total bits, below which degradation becomes apparent.
+I therefore selected ap_fixed<13,5> (13 total bits, 5 integer bits) as the default precision for FPGA synthesis.
+5.2 Float vs. Quantised Model Performance
+The effect of quantisation on classification performance was evaluated by comparing the floating-point and ap_fixed<13,5> implementations on the same validation datasets.
+As shown in Figure 6b, the ROC curves for the two models are nearly identical, with the AUC decreasing only marginally from the PyTorch floating-point baseline () to the quantised implementation ().
+This confirms that the chosen fixed-point precision preserves the model’s separation power between signal and background, enabling an efficient FPGA implementation without loss of physics performance.
+5.3 FPGA Synthesis with Vitis 2023
+The quantised model was synthesised using Vitis HLS 2023.2 [28], targeting three representative FPGA families: the low-cost Zynq-7000 (xc7z020), the mid-range Kintex UltraScale (xcku040), and the high-end Alveo U250 (xcu250).
+A target clock period of () was assumed in all cases.
+The measured end-to-end latency, pipeline initiation interval (II), and DSP usage are summarised in Table 1.
+| Platform | Latency [s] | II | DSP usage | Timing met |
+| Zynq-7000 xc7z020 | 6.51 | 1060 | 963 % | yes |
+| Kintex UltraScale xcku040 | 3.30 | 654 | 108 % | yes |
+| Alveo U250 xcu250 | 1.99 | 394 | 17 % | yes |
+5.3.1 Resource Utilisation
+DSP slices are the critical limiting resource for this design, as they are heavily used for multiply–accumulate (MAC) operations in convolutional and dense layers.
+While LUT and BRAM usage remain well below device limits in all cases, DSP saturation is encountered for the Zynq-7000 and, to a lesser extent, the Kintex UltraScale.
+This constraint dictates optimisation strategies, such as resource sharing or further model compression, for deployment on cost-sensitive platforms.
+The Alveo U250, by contrast, offers substantial headroom for scaling up the model or reducing latency further.
+5.3.2 Latency
+Figure 7 shows the measured inference latency for the three synthesis targets.
+Even without aggressive optimisation, the high-end Alveo U250 achieves sub- latency.
+The Kintex UltraScale operates at , while the Zynq-7000 delivers .
+All results meet the -level latency guideline for station-level triggering in the target application.
+5.4 Power consumption
+I characterised fabric-only on-chip power for the HLS4ML-generated accelerator implemented on a Xilinx Alveo U250 (xcu250-figd2104-2L-e) using Vivado 2023.2.
+To measure only the neural-network accelerator’s contribution, I built a minimal ‘fabric-only’ design: the HLS4ML core was the sole processing block at the top level, with no platform shell, memory controllers, PCIe or other peripherals, and no external I/O interfaces.
+The only additional logic was an on-chip pseudo-random bit sequence (PRBS) generator to provide a continuous input stimulus.
+The design was driven by a single clock and taken through full implementation (synthesis, placement, and routing) so power was reported on the placed-and-routed netlist.
+The PRBS source was implemented as a 32-bit LFSR with taps , replicated across the input bus, with handshake/control pins held active to keep the pipeline filled.
+This stimulus exercises the full datapath and prevents constant-propagation through the network.
+Post-implementation power was obtained with Vivado’s vectorless analysis (no activity annotation).
+The placed-and-routed netlist reports a dynamic power of for the core.
+If clock-network power is also excluded, the dynamic component attributable solely to datapath logic is approximately .
+Because the estimate remains vectorless and excludes platform infrastructure, the reported numbers should be interpreted as a lower bound; absolute power is expected to increase with activity-annotated runs and board-level I/O.
+As the U250 is the largest of the three target platforms, this value is conservative with respect to device scaling.
+These results indicate ample headroom within a typical budget for a compact radio-detection pipeline, leaving margin for front-end filtering, data conversion, and control logic.
+A full platform-inclusive power profile (including activity-annotated power and comparisons of LUT- vs.
+DSP-mapped arithmetic) is left for future hardware studies.
+6 Summary and Outlook
+I have presented an artificial-intelligence-based self-trigger for radio detection of extensive air showers and demonstrated its feasibility for real-time deployment on FPGA hardware.
+Starting from a floating-point model trained in PyTorch on measured noise from a high-interference environment combined with cosmic-ray pulses generated using the MGMR3D simulation framework, the workflow covers a broad range of shower geometries, energies, and antenna–core distances.
+The trained fully convolutional classifier achieves stable performance under operating conditions representative of MHz-scale trigger-trial rates encountered in urban or RFI-rich sites.
+For context, I also compared the neural-network classifier to a simple threshold-based trigger, implemented as an SNR cut on the Hilbert-envelope peak relative to the noise variance.
+The baseline achieves only modest separation power () and a signal efficiency of at the target operating point of , whereas the classifier retains efficiency under the same conditions.
+This comparison demonstrates that the AI-based approach provides a substantial improvement over traditional threshold triggers in the challenging high-interference environment considered here.
+The network was quantised using HLS4ML to a fixed-point precision of ap_fixed<13,5>.
+A precision scan confirmed that 13-bit fixed-point arithmetic preserves the physics performance of the floating-point model (, ).
+The quantised model was synthesised with Vitis HLS 2023.2 for three FPGA platforms (Zynq-7000, Kintex UltraScale, and Alveo U250).
+High-level synthesis results show inference latencies in the microsecond range, meeting the first-level trigger timing budget for large-scale radio arrays.
+DSP usage emerged as the primary resource bottleneck, guiding optimisation strategies for deployment on cost-constrained devices.
+A post-implementation, fabric-only power estimate on the Alveo U250 reports at (vectorless analysis), which I interpret as a conservative lower bound on on-chip power.
+For the datasets discussed in this work, strong narrow-band RFI components were not removed prior to training.
+In a real-world deployment, eliminating such RFI using an FFT-based filter implemented on the FPGA could simplify the classification task by reducing background complexity.
+A simpler network architecture may lead to faster inference times and lower computational costs, paving the way for real-time triggering with further reduced latency and resource usage.
+This work demonstrates, for the first time in the context of radio-detector self-triggering, an end-to-end workflow from AI model training to FPGA-oriented deployment using HLS4ML: quantisation to fixed-point, high-level synthesis with Vitis HLS 2023.2, and validation of physics-performance preservation together with post-synthesis latency/resource characterisation and a fabric-only, post-implementation power estimate.
+The approach reduces reliance on external triggers from particle or fluorescence detectors and supports autonomous operation scenarios for next-generation radio arrays.
+Future developments include broadening the training to a wider family of simulated pulse morphologies and incorporating additional measured air-shower events.
+On the hardware side, it will be important to complement the fabric-only power estimate reported here with activity-annotated and platform-inclusive profiles across the target devices, and to investigate architectural and clocking optimisations suitable for low-power autonomous stations.
+Together with simple on-FPGA pre-filtering of narrow-band RFI, these directions are expected to further reduce resource usage and latency and to de-risk deployment in operational radio arrays.
+In addition to these technical aspects, the comparison with a simple SNR-threshold baseline highlights the physics impact of the proposed method.
+At the chosen operating point, the neural-network trigger lowers the effective trigger threshold significantly compared to the SNR baseline.
+This improved threshold increases sensitivity to weak signals and to very inclined air showers, where the particle cascade is absorbed while the radio emission remains detectable.
+Such events are essential not only for composition studies and for extending the sky coverage of cosmic-ray observatories, but also as a prerequisite for detecting ultra-high-energy neutrinos with ground-based radio arrays.
+Moreover, the approach is designed to operate robustly in environments with strong and variable radio backgrounds, where conventional threshold triggers fail, thereby enabling autonomous operation under realistic field conditions.
+These results demonstrate that FPGA-based AI self-triggering not only meets the technical feasibility requirements but also provides a tangible enhancement of the scientific potential of next-generation radio and neutrino observatories.
 Acknowledgments
@@
 Their contributions facilitated the development and testing of the hardware components, streamlining the experimental setup.
-This research made use of open-source software, including PyTorch and various scientific computing libraries in Python, such as NumPy, SciPy, and Matplotlib.
+The author also gratefully acknowledges Prof.
+Katie Mulrey and Prof.
+Harm Schoorlemmer (Radboud University Nijmegen) for kindly providing the antenna employed in this study.
+The assistance of Noah Siegemund in setting up the hardware on campus is also gratefully acknowledged.
+This research made use of open-source software, including PyTorch, HLS4ML and various scientific computing libraries in Python, such as NumPy, SciPy, and Matplotlib.
 The author acknowledges their developers for providing these valuable tools, which played a crucial role in data processing, model training, and analysis.